OpenStudy (anonymous):

Solve. 2x2 + 9x = 35

OpenStudy (anonymous):

\[2x^2+9x-35=0\] maybe it factors, if not use teh quadratic formula

OpenStudy (anonymous):

try \((x+7)(2x-5)=0\)

OpenStudy (anonymous):

I did that but i stopped at 14x and -5x to equal tthe 9

OpenStudy (anonymous):

factoring is the key to this one once you have it factored, solve \[x+7=0\] and \[2x-5=0\]

OpenStudy (anonymous):

ohhh okay thanks :D